OpenStudy (anonymous):

linear equation: x/2-1=5/6

sam (.sam.):

(x)/(2)-1=(5)/(6) x/2=11/6 x=11/3

OpenStudy (anonymous):

where did the 11 come from??

sam (.sam.):

\[\frac{5}{6}+1\] \[\frac{5}{6}+\frac{6}{6}\] \[\frac{5+6}{6}\] \[\frac{11}{6}\]

OpenStudy (anonymous):

oohhh ok thank u